OpenStudy (unklerhaukus):

\[(2t^3y-y^3)\text dt-2t^4\text dy=0\]

OpenStudy (unklerhaukus):

\[\frac{\partial M}{\partial y}=2t^3-3y^2\]\[\frac{\partial N}{\partial t}=-8t^3\]\[\frac{\partial M}{\partial y}\neq\frac{\partial N}{\partial t}\]

OpenStudy (unklerhaukus):

\[R=R(t,y)=t^my^m\]\[\frac{\partial MR(t,y)}{\partial y}=\frac{\partial NR(t,y)}{\partial t}\]\[R\frac{\partial M}{\partial y}+R_yM=R\frac{\partial N}{\partial t}+R_tN\]

OpenStudy (unklerhaukus):

\[R\left(\frac{\partial M}{\partial y}-\frac{\partial N}{\partial t}\right)=R_tN-R_yM\] \[t^my^n\left(2t^3y-3y^2+8t^3\right)=mt^{m-1}y^n(2t^3y−y^3)+nt^my^{n-1}(2t^4)\]

OpenStudy (unklerhaukus):

*\[R=R(t,y)=t^my^n\]

OpenStudy (unklerhaukus):

how to find \(m,n\)

OpenStudy (anonymous):

I haven't tried this yet, but you're trying to solve a exact equation, mind if I ask you what you did to make this exact? There is this annoying negative sign

OpenStudy (unklerhaukus):

i assumed an intergrating facor was of the form \( R(t,y)=t^my^n\)

OpenStudy (unklerhaukus):

which negative sign do you mean?

OpenStudy (dominusscholae):

It seems the method of homogenous equations is an easier method of solving this problem.

OpenStudy (unklerhaukus):

well im trying to do it this way

OpenStudy (anonymous):

\[M(x.y)+N(x,y)\] hmm yes I understand that part, what bothers me is that the function above doesn't look like a characteristcal exact equation to me. Sorry it's a pain to LaTeX here at the moment, I am on my DOs machine, not used to it.

OpenStudy (anonymous):

If I recall it correctly, the plus in between is essential to be exact.

OpenStudy (unklerhaukus):

N=-N ?

OpenStudy (dominusscholae):

Ok I solved it. Only thing is I didn't use an integrating factor unfortunately. Lemme try to see if I can.

OpenStudy (anonymous):

I will have to work through it myself at first too, sorry if I caused more questions then answers, but that negative sign really seems alien to me at first glance, various substitution will sure get rid of it, I just can't recall if that influences the result or not.

OpenStudy (unklerhaukus):

am i going about this the right way?

OpenStudy (dominusscholae):

To be honest, it doesn't seem like it. The most feasible thing to find m and n in this case is to solve\[R(dM/dy - dN/dt) = 0\] but it doesn't look like you can get a solution by hand. Since this is a homogenous equation, I suggest you tackle that direction since it's a lot more straightforward.

OpenStudy (unklerhaukus):

\[R\frac{\partial M}{\partial y}+R_yM=R\frac{\partial N}{\partial t}+R_tN\] \[t^my^n(2t^3)+nt^my^{n-1}(2t^3y)=t^my^n(-8t^3)+mt^{m-1}y^n(-2t^4)\] \[2t^{m+3}y^n+2nt^{m+3}y^{n}=-8t^{m+3}y^n-2mt^{m+3}y^n\] \[(10+2n+2m)t^{m+3}y^n=0\] \[(10+2n+2m)=0\]

OpenStudy (dominusscholae):

Ahh yes, sorry I meant to say solve for \[d/dy(RM) = d/dx(RN)\]

OpenStudy (unklerhaukus):

im not sure what your saying @dominusscholae

OpenStudy (dominusscholae):

^basically what you did in the last post. You seem to be doing the correct thing in order to find m and n. So (at least to me) any combination of m and n that satisfies 10 +2m + 2n = 0. Thing is, when I plug any value satisfying this, I'm not getting an exact equation for some reason.

OpenStudy (unklerhaukus):

i need two equations to solve two variables, how do i get the other/